# Peng Cheng, and Jiming Chen

### v2

@@ Abstract
Reinforcement learning (RL) is a powerful tool for sequential decision-making, but its application is often hindered by privacy concerns arising from its interaction data. This challenge is particularly acute in advanced Cyber-Physical Systems (CPS), where learning from operational and user data can expose systems to privacy inference attacks. Existing differential privacy (DP) models for RL are often inadequate: the centralized model requires a fully trusted server, creating a single point of failure risk, while the local model incurs significant performance degradation that is unsuitable for many control applications. This paper addresses this gap by leveraging the emerging shuffle model of privacy, an intermediate trust model that provides strong privacy guarantees without a centralized trust assumption. We present Shuffle Differentially Private Policy Elimination (SDP-PE), the first generic policy elimination-based algorithm for episodic RL under the shuffle model. Our method introduces a novel exponential batching schedule and a “forgetting” mechanism to balance the competing demands of privacy and learning performance. Our analysis shows that SDP-PE achieves a near-optimal regret bound, demonstrating a superior privacy-regret trade-off that significantly outperforms the local model. This work establishes the viability of the shuffle model for secure data-driven control in advanced CPS.
Index Terms:
Reinforcement Learning, Shuffle Differential Privacy, Online Learning.
@@ I Introduction
Reinforcement learning (RL) is a sequential decision-making framework that has gained remarkable attraction for CPS, such as smart grid, networked robotics, and autonomous vehicle networks. In these systems, a learning agent continuously interacts with the environment to receive new observations and rewards, and improves the performance of its policy. However, a crucial concern for its deployment is that the interaction data usually contains sensitive information. This concern is not only limited to personal data in user-centric CPS [1], but also extends to confidential operational data, such as control parameters or system states in a distributed network [2]. Without privacy protection mechanisms in place, the learning agent can memorize sensitive information from the interaction history [3], exposing confidential operational patterns or leaving the system’s underlying entities vulnerable to various privacy attack [4].
To counter such threat, differential privacy (DP) [5] has been extensively applied as a formal privacy method in various decision-making settings, including multi-armed bandits [6], classical control systems [7, 8], and networked control systems [9, 10]. In the case of RL, a growing body of literature has addressed the privacy of different components of the learning process, such as protecting the reward function [11], the transition dynamics [12], or the learned policies [13]. Distinct from this line of work, our focus is on protecting the interaction history itself. Prior work has primarily focused on two privacy architectures: the centralized privacy model with Joint DP (JDP) constraint [14, 15] and the local privacy model with Local DP (LDP) constraint [16, 17]. To apply these frameworks, the standard approach is to consider an episodic online interaction setting. In this formulation, the agent interacts with the environment for $K$ episodes, and each episode is regarded as an interaction with one specific user. In a CPS context, this “user” could be a human customer, a smart home, or an autonomous vehicle providing data. The aim is to ensure that the user’s states and generated rewards will not be inferred by an adversary during the learning process. Specifically, JDP guarantees that all other users’ decisions will not leak much information about any specific user, and $\varepsilon$-JDP can be obtained at the expense of an additive logarithmic term of $\mathcal{O}(\log(K)/\varepsilon)$ in the regret bound [14]. However, a trusted central server is required to collect and process the raw interaction data under JDP, which may result in a single point of failure. On the other hand, $\varepsilon$-LDP removes this risk by privatizing data on each users’ local device, but at the cost of severe performance degradation, resulting in a multiplicative regret term $\mathcal{O}(\sqrt{K}/\varepsilon)$ [16] that is often unacceptable for control applications. This naturally leads to the following question: Can a finer trade-off between privacy and regret in RL be achieved, i.e., achieving utility comparable to that of centralized privacy models but without relying on a trusted central server?
@@
Motivated by these, this paper leverages an intermediate trust model of differential privacy, known as shuffle differential privacy (SDP) [18], as a resilient, distributed architecture for private RL. As illustrated in Figure 1, this paradigm uses a secure shuffler to anonymize a batch of locally randomized data before it reaches the central server, and makes sure that the central server can not distinguish any two users’ data. The shuffler is trusted by these users since it can be implemented by trusted cryptographic mixnets or trusted hardware deployed via distributed constructs [19]. This privacy model provides a stronger privacy guarantee than the centralized model but usually suffers a smaller cost than the local model. The key intuition is that additional randomness of the shuffler creates a privacy blanket [20] so that each user can inject much less noise and hide their information in the crowd. A better privacy-utility trade-off has been proven in related fields like (federated) supervised learning [21, 22], and bandits [23, 24]. However, to the best of our knowledge, the application of the shuffle model in RL remains largely unexplored. This gap is due to a fundamental tension between the batch-processing requirement of shuffle privacy and the sequential, online nature of most RL algorithms. Specifically, the shuffle model introduces two compounding challenges. First, by operating on anonymized batches, it imposes an inherent delay, as the agent receives no feedback until an entire batch of interactions is complete. Second, to ensure privacy, the aggregated data is intentionally perturbed with statistical noise. This combination of delayed feedback and noisy observations severely complicates the crucial exploration-exploitation trade-off, posing a significant challenge to developing private RL algorithms that can achieve near-optimal regret while satisfying the SDP constraint.
The batch-based updates required by SDP makes the literature on multi-batch RL highly relevant. This field aims to improve deployment efficiency by minimizing the number of policy switches while maintaining low regret. Approaches in this area are typically either adaptive [25, 26], where batch sizes are determined adaptively based on online data collection, or static [27, 28], where the batch schedule is fixed in advance and offer no intra-batch adaptivity. This line of work has produced algorithms that limit the number of policy switches to an optimal $\mathcal{O}(\log\log K)$, while maintaining an optimal $\tilde{\mathcal{O}}(\sqrt{K})$ regret. The static approach is more relevant to our context, as its core constraint, not using data until a batch is complete, naturally aligns with the shuffle privacy protocol. However, these existing methods are not directly applicable since they can access accurate interaction histories to optimize the policy. In contrast, our algorithm must be robust to the significant noise injected to ensure privacy. This requires a fundamentally different design, one where the batching mechanism and behavior policies are carefully co-designed to control and mitigate the impact of these privacy-induced perturbations.
@@
In this paper, we overcome these challenges by introducing Shuffle Differentially Private Policy Elimination (SDP-PE), the first RL algorithm designed to satisfy the constraint of shuffle privacy. We provide both rigorous theoretical analysis and empirical validation for our method, and the main contributions are threefold:
@@
  • Generic Algorithm Design: Based on the policy elimination idea [27], we develop two key strategies tailored for the shuffle model: a) a staged learning process with exponentially increasing batch sizes. This structure allows for the iterative refinement of an active policy set, and is critical for managing the exploration-exploitation trade-off under the delayed feedback inherent to the shuffle model. b) a “forgetting” mechanism that uses only data from the current stage for model updates. This approach prevents the accumulation of privacy-induced noise across the learning stages.
@@
DP-UCBVI  [29] $\varepsilon$-JDP $\widetilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+\frac{X^{2}AH^{3}}{\varepsilon}\right)$
@@
The remainder of this paper is organized as follows. In Section II, we provide the preliminaries of episodic RL and adapt Shuffle DP for RL and formulate the problem. Section III presents the designed policy elimination algorithm under SDP constraints. The regret and privacy guarantees are presented in Sections IV and V. The numerical experiment is shown in Section VI. Discussion and conclusion are presented in Section VII and VIII, respectively.
@@ II-A Episodic Reinforcement Learning
We consider an agent (e.g., a controller) tasked with learning to interact with unknown dynamics, modeled as an MDP. The agent interacts with the environment for $K$ episodes. In the security-conscious CPS setting, each episode $k\in[K]$ is modeled as an interaction with a unique system component or user $u_{k}\in\mathcal{U}$, where $\mathcal{U}$ is the user space. For each episode $k\in[K]$, the learning agent determines policy $\pi_{k}$ and sends it to user $u_{k}$ for execution. The output of the execution, a trajectory $S_{k}=\left(x_{h}^{k},a_{h}^{k},r_{h}^{k}\right)_{h\in[H]}$ representing the sensitive data, is sent back to the learning agent for updating the policy. We measure the performance of a learning algorithm by its cumulative regret after $K$ episodes,
$\text{Regret}(K):=\sum_{k=1}^{K}\left[V_{1}^{\star}(x_{1}^{k})-V_{1}^{\pi_{k}}(x_{1}^{k})\right],$ (1)
where $x_{1}^{k}$ is the initial state sampled from $d_{1}$.
@@ II-B Differential Privacy and Shuffle Model
Differential Privacy provides a formal guarantee of data privacy. Consider a general case where $\mathcal{S}$ is the data universe, and we have $n$ unique users. We say $D,D^{\prime}\in\mathcal{S}^{n}$ are neighboring batched datasets if they only differ in one user’s data for some $i\in[n]$. Then, we have the standard definition of DP [5]:
@@
The shuffle model is an enhanced DP protocol $\mathcal{T}=(\mathcal{E},\mathcal{F},\mathcal{G})$ involving three parts, as shown in Figure 1: (i) a (local) random encoder $\mathcal{E}$ at each user’s side; (ii) a secure shuffler $\mathcal{F}$; and (iii) an analyzer $\mathcal{G}$ at the central server (i.e., the learning agent). In this model, each user $u_{i}$ first locally randomizes their data $D_{i}$ via the encoder $\mathcal{E}$, and sends the resulting messages $\mathcal{E}(D_{i})$ to the shuffler $\mathcal{F}$. The shuffler $\mathcal{F}$ permutes messages from this batch users at random and then reports $\mathcal{F}(\mathcal{E}(D_{1}),\dots,\mathcal{E}(D_{n}))$ to the analyzer. The analyzer $\mathcal{G}$ then aggregates the received messages and outputs desired statistics. In this protocol, the users trust the shuffler but not the analyzer, and the goal is to ensure the output of the shuffler $\mathcal{F}$ on two neighboring datasets is indistinguishable in the analyzer’s view. Here, we define the mechanism $(\mathcal{F}\circ\mathcal{E}^{n})(D)=\mathcal{F}(\mathcal{E}(D_{1}),\dots,\mathcal{E}(D_{n}))$, where $D\in\mathcal{D}^{n}$. Thus, we have the definition of standard shuffle DP [18]:
@@
In a sequential data-generation setting, this protocol will be applied across multiple batches. To protect the sequence of outputs across all $B$ batches, we define the (composite) mechanism $\mathcal{M}_{\mathcal{T}}=(\mathcal{F}\circ\mathcal{E}^{L_{1}},\dots,\mathcal{F}\circ\mathcal{E}^{L_{B}}),$ where each mechanism $\mathcal{F}\circ\mathcal{E}^{L_{b}}$ operates on a dataset of $L_{b}$ users. This leads to the following definition $B$-batch SDP [18]:
@@ II-C Problem Formulation: Shuffle-Private RL
Here, we formally define the private reinforcement learning problem. We consider an agent interacting with an unknown episodic MDP as described in Section II-A, but under the communication constraint of the $B$-batch shuffle privacy defined in Section II-B.
@@ III Algorithm Design
In this section, we introduce a generic algorithmic framework, Shuffle Differentially Private Policy Elimination (SDP-PE, Algorithm 1). Unlike existing private RL algorithms that rely on the optimism in the face of uncertainty principle [14, 16], SDP-PE builds on the policy elimination idea from [27, 28], which extends action elimination from bandits [30] to RL.
@@
The high-level, three-step structure of our algorithm is inspired by the APEVE framework of [27]. However, our objective of minimizing regret under shuffle privacy, rather than minimizing policy switches, necessitates several fundamental innovations in the design as detailed below.
@@
Figure 2: A visualization of the procedures of SDP-PE (Algorithm 1).
@@ III-B Crude Exploration in Algorithm 2
This set is then used to construct an absorbing MDP, a technique adapted from [27]. As fomally defined below, this method isolates the hard-to-reach states by redirecting their transition probabilities to a single absorbing state, $x^{\dagger}$.
@@ III-C Fine Exploration in Algorithm 3
This policy is specifically designed to minimize the “worst-case coverage number” over the current set of active policies, which could provide uniform coverage of all remaining policies. To generate data, the users execute both this new policy $\pi^{\text{ref},b}$ and the auxiliary policy $\pi_{0}$ from crude exploration phase. The combined dataset $D$ is then processed by the Privatizer to yield a refined private estimate $\widetilde{P}^{\text{ref},b}$ and reward estimate $\widetilde{r}^{b}$, which are more accurate than their crude counterparts. Notably, different from the APEVE algorithm [27], our fine exploration policy is specifically designed to achieve better policy coverage, leading to tighter confidence bounds and an improved final regret.
The final step is each stage is policy elimination. We evaluate all policies in $\phi_{b}$ using the reliable estimates $\widetilde{P}^{\text{ref},b}$ and $\widetilde{r}^{b}$, and update the active policy set by eliminating all policies whose upper confidence bound (UCB) is less than the lower confidence bound (LCB) of any other active policy. The confidence intervals shrink along with the policy elimination step, ensuring that the optimal policy stays in the active policy set with high probability, which guarantees that the algorithm eventually converges to the optimal policy.
IV Regret Guarantee
The main theoretical result of this paper is the following regret bound for our SDP-PE algorithm.
Theorem 6 (Regret bound of SDP-PE)

### v3

@@ Abstract
Reinforcement learning (RL) is a powerful tool for sequential decision-making, but its application is often hindered by privacy concerns arising from its interaction data. This challenge is particularly acute in advanced networked systems, where learning from operational and user data can expose systems to privacy inference attacks. Existing differential privacy (DP) models for RL are often inadequate: the centralized model requires a fully trusted server, creating a single point of failure risk, while the local model incurs significant performance degradation that is unsuitable for many networked applications. This paper addresses this gap by leveraging the emerging shuffle model of privacy, an intermediate trust model that provides strong privacy guarantees without a centralized trust assumption. We present Shuffle Differentially Private Policy Elimination (SDP-PE), the first generic policy elimination-based algorithm for episodic RL under the shuffle model. Our method introduces a novel exponential batching schedule and a “forgetting” mechanism to balance the competing demands of privacy and learning performance. Our analysis shows that SDP-PE achieves a near-optimal regret bound, demonstrating a superior privacy-regret trade-off with utility comparable to the centralized model while significantly outperforming the local model. The numerical experiments also corroborate our theoretical results and demonstrate the effectiveness of SDP-PE. This work establishes the viability of the shuffle model for secure data-driven decision-making in networked systems.
@@ I Introduction
Reinforcement learning (RL) has gained remarkable attraction to optimize networked systems [8714026], such as wireless network, smart grid, and autonomous vehicle networks. A common operational paradigm is online reinforcement learning, as illustrated in Figure 1: a central agent (e.g., a network coordinator) performing ”centralized learning” to optimize a global policy, while providing streaming services to a sequence of online nodes or user requests. For each request, the central agent deploys its global policy to the corresponding distributed node (e.g., 5G small cell or cognitive radio node). The node locally and autonomously executes this policy to generate the interaction trajectory (e.g., an $H$-steps communication session), and the trajectory is then sent back to the central agent for updating the global policy. For instance, in Dynamic Spectrum Access (DSA), a cognitive radio node runs the policy locally, sensing its local spectrum state to choose a channel. The coordinator’s goal is to learn a global operational policy that optimizes network-wide performance (e.g., maximizing spectral efficiency) using the interaction data fed back by the nodes.
Figure 1: An illustration of online reinforcement learning in a distributed network. A central coordinator performs ”centralized learning”, and provides service for online streaming user requests in distributed nodes by policy deployment. Each local node generates episodic interaction data. This sensitive interaction data is then (privately) transmitted back to the coordinator for global policy updates.
However, a crucial concern for its deployment is that the interaction data from each node usually contains sensitive information (e.g., local traffic patterns and spectrum usage habits in the DSA instance). Without privacy protection mechanisms in place, the learning agent can memorize sensitive information from the interaction history [zhang2016privacy], leaving the system to various privacy attack [mo2024security]. Existing Differential Privacy (DP) [dwork2014algorithmic] models for such sequential decision-making [vietri2020private, baidifferentially, garcelon2021local, chowdhury2022differentially] are often inadequate: the centralized privacy model (JDP) [vietri2020private, baidifferentially] requires a fully trusted server to collect raw trajectories, creating a single point of failure. Conversely, the local privacy model (LDP) [garcelon2021local, chowdhury2022differentially] avoids this risk by privatizing data on each node’s device, but at the cost of severe performance degradation (e.g., a multiplicative regret term $\mathcal{O}(\sqrt{K}/\epsilon)$), which is often unacceptable for performance-critical network optimization. This naturally leads to the following question: Can a finer trade-off between privacy and regret in RL be achieved, i.e., achieving utility comparable to that of centralized privacy models but without relying on a trusted central server?
@@
Motivated by these, this paper leverages an intermediate trust model of differential privacy, known as shuffle differential privacy (SDP) [cheu2019distributed], as a resilient, distributed architecture. As illustrated in Figure 2, this paradigm requires a secure shuffler to anonymize a batch of locally randomized data before it reaches the central server, and makes sure that the central server can not distinguish any two users’ data. The shuffler, implemented by trusted cryptographic mixnets [bittau2017prochlo], creates a “privacy blanket” [balle2019privacy] so that each user can inject much less noise and hide their information in the crowd. This allows each node to inject significantly less noise than in LDP, enabling the system to achieve high utility comparable to JDP but without the centralized trust assumption, making it highly practical.
However, the batch-processing requirement of SDP forces us to abandon the classical “per-interaction” online model and shifts to a “batched” online model. This privacy-driven “batching” constraint poses new algorithmic challenges. First, by operating on anonymized batches, it imposes an inherent delay, as the agent receives no feedback until an entire batch of interactions is complete. Second, to ensure privacy, the aggregated data is intentionally perturbed with statistical noise. This combination of delayed feedback and noisy observations severely complicates the crucial exploration-exploitation trade-off. At the same time, this constraint also creates an opportunity to solve a second critical engineering bottleneck in this streaming service model: the policy switching cost. Standard online algorithms require per-interaction “reconfiguration” updates (i.e., after every $H$-step session), which is operationally infeasible in large-scale networks due to the heavy deployment and communication cost [8500100]. The “batched” model, necessitated by SDP, naturally allows for batched policy updates, providing a path to solve this engineering problem.
@@
In this paper, we overcome these challenges by introducing Shuffle Differentially Private Policy Elimination (SDP-PE), the first online RL framework co-designed to exploit this synergy. It uses an exponential schedule that both satisfies the batching needs of SDP and solves the engineering cost of policy switching. We provide both rigorous theoretical analysis and empirical validations for our method, and the main contributions are summarized as follows:
  • Problem Formulation: We formalize the “distributed-to-global” network optimization problem for streaming online services (like DSA) under the dual considerations of SDP and low policy switching cost.
@@
  • Algorithm Design: Based on the policy elimination idea [qiao2022sample], we develop two key strategies tailored for the shuffle model: a) a staged learning process with exponentially increasing batch sizes. This structure allows for the iterative refinement of an active policy set, and is critical for managing the exploration-exploitation trade-off under the delayed feedback inherent to the shuffle model. b) a “forgetting” mechanism that uses only data from the current stage for model updates. This approach prevents the accumulation of privacy-induced noise across the learning stages.
@@
  • Numerical Validations: We empirically demonstrate our framework’s (a) superior privacy-utility trade-off, (b) massive (orders of magnitude) reduction in network switching/deployment costs, which also further corroborates our theoretical findings.
@@
DP-UCBVI  [qiao2023near] $\varepsilon$-JDP $\widetilde{\mathcal{O}}\left(\sqrt{XAH^{3}K}+\frac{X^{2}AH^{3}}{\varepsilon}\right)$
@@
The remainder of this paper is organized as follows. In Section II, we provide the preliminaries of episodic RL and adapt Shuffle DP for RL and formulate the problem. Section III presents the designed policy elimination algorithm under SDP constraints. The regret and privacy guarantees are presented in Sections IV and V. The numerical experiment is shown in Section VI. Discussion and related work are presented in Section VII and Section VIII, respectively. Finally, Section IX presents the conclusion.
@@ II-A Episodic Reinforcement Learning
We consider a central agent tasked with interacting with a sequence of $K$ online user requests in the network system. Each episode $k\in[K]$ is modeled as an interaction with user request $u_{k}\in\mathcal{U}$, where $\mathcal{U}$ is the user space. For each episode $k\in[K]$, the learning agent determines its current policy $\pi_{k}$ and deploys it to user $u_{k}$ for execution. The output of the execution, a trajectory $S_{k}=\left(x_{h}^{k},a_{h}^{k},r_{h}^{k}\right)_{h\in[H]}$, represents the sensitive data to be protected. The trajectory is then sent back to the learning agent for policy updating. We measure the performance of a learning algorithm by its cumulative regret after $K$ episodes,
$\text{Regret}(K):=\sum_{k=1}^{K}\left[V_{1}^{\star}(x_{1}^{k})-V_{1}^{\pi_{k}}(x_{1}^{k})\right],\quad x_{1}^{k}\sim d_{1}.$ (1)
A secondary metric is the policy switching cost [bai2019provably], which measures how many times the agent changes the policy,
$N_{switch}:=\sum_{k=1}^{K-1}\mathbf{1}\{\pi_{k}\neq\pi_{k+1}\}.$ (2)
@@ II-B Differential Privacy and Shuffle Model
Differential privacy provides a formal guarantee of data privacy. Consider a general case where $\mathcal{S}$ is the data universe, and we have $n$ unique users. We say $D,D^{\prime}\in\mathcal{S}^{n}$ are neighboring batched datasets if they only differ in one user’s data for some $i\in[n]$. Then, we have the standard definition of DP [dwork2014algorithmic]:
@@
The shuffle model is an enhanced DP protocol $\mathcal{T}=(\mathcal{E},\mathcal{F},\mathcal{G})$ involving three components, as shown in Figure 2: (i) a (local) random encoder $\mathcal{E}$ at each user’s side; (ii) a secure shuffler $\mathcal{F}$; and (iii) an analyzer $\mathcal{G}$ at the central server (i.e., the learning agent). In this model, each user $u_{i}$ first locally randomizes their data $D_{i}$ via the encoder $\mathcal{E}$, and sends the resulting messages $\mathcal{E}(D_{i})$ to the shuffler $\mathcal{F}$. The shuffler $\mathcal{F}$ permutes messages from this batch users at random and then reports $\mathcal{F}(\mathcal{E}(D_{1}),\dots,\mathcal{E}(D_{n}))$ to the analyzer. The analyzer $\mathcal{G}$ then aggregates the received messages and outputs desired statistics. In this protocol, the users trust the shuffler but not the analyzer, and the goal is to ensure the output of the shuffler $\mathcal{F}$ on two neighboring datasets is indistinguishable in the analyzer’s view. Here, we define the mechanism $(\mathcal{F}\circ\mathcal{E}^{n})(D)=\mathcal{F}(\mathcal{E}(D_{1}),\dots,\mathcal{E}(D_{n}))$, where $D\in\mathcal{D}^{n}$. Thus, we have the definition of standard shuffle DP [cheu2019distributed]:
@@
In a sequential data-generation setting, this protocol will be applied across multiple batches. To protect the sequence of outputs across all $B$ batches, we define the (composite) mechanism $\mathcal{M}_{\mathcal{T}}=(\mathcal{F}\circ\mathcal{E}^{L_{1}},\dots,\mathcal{F}\circ\mathcal{E}^{L_{B}}),$ where each mechanism $\mathcal{F}\circ\mathcal{E}^{L_{b}}$ operates on a dataset of $L_{b}$ users. This leads to the following definition $B$-batch SDP [cheu2019distributed]:
@@ II-C Problem Formulation: Shuffle-Private RL
Building upon the above definitions, we formally define the problem addressed in this paper. We consider an agent interacting with an unknown episodic MDP as described in Section II-A, but under the communication constraint of the $B$-batch shuffle privacy defined in Section II-B.
@@ III Algorithm Design
In this section, we introduce a generic algorithmic framework, Shuffle Differentially Private Policy Elimination (SDP-PE, Algorithm 1). Unlike existing private RL algorithms that rely on the optimism in the face of uncertainty principle [vietri2020private, garcelon2021local], SDP-PE builds on the policy elimination idea from [qiao2022sample, zhang2022near], extending action elimination from bandits [evendar06a] to RL.
@@
The high-level, three-step structure of our algorithm is inspired by the APEVE framework of [qiao2022sample]. However, our objective of minimizing regret under shuffle privacy, rather than minimizing policy switches, necessitates several fundamental innovations in the design as detailed below.
@@
Figure 3: Overview of the SDP-PE algorithm (Algorithm 1). The central coordinator deploys a global policy to distributed nodes, which process batched user requests. Private interaction data is then shuffled and transmitted back for policy value updates and elimination.
@@ III-B Crude Exploration in Algorithm 2
This set is then used to construct an absorbing MDP, a technique adapted from [qiao2022sample]. As fomally defined below, this method isolates the hard-to-reach states by redirecting their transition probabilities to a single absorbing state, $x^{\dagger}$.
@@ III-C Fine Exploration in Algorithm 3
This policy is specifically designed to minimize the “worst-case coverage number” over the current set of active policies, which could provide uniform coverage of all remaining policies. To generate data, the users execute both this new policy $\pi^{\text{ref},b}$ and the auxiliary policy $\pi_{0}$ from crude exploration phase. The combined dataset $D$ is then processed by the Privatizer to yield a refined private estimate $\widetilde{P}^{\text{ref},b}$ and reward estimate $\widetilde{r}^{b}$, which are more accurate than their crude counterparts. Notably, different from the APEVE algorithm [qiao2022sample], our fine exploration policy is specifically designed to achieve better policy coverage. This design choice guarantees a more uniform value estimation accuracy for all candidate policies, leading to tighter confidence bounds and an improved final regret bound.
The final step of each stage is policy elimination, as illustrated in Figure 4. We evaluate all active policies in $\phi_{b}$ using the reliable estimates $\widetilde{P}^{\text{ref},b}$ and $\widetilde{r}^{b}$, and update the active policy set by eliminating all policies whose upper confidence bound (UCB) falls below the lower confidence bound (LCB) of any other active policy. The confidence intervals shrink along with the policy elimination step, ensuring that the optimal policy stays in the active policy set with high probability, which guarantees that the algorithm eventually converges to the optimal policy.
Figure 4: Example of policy elimination. The LCB of the estimated optimal policy $\pi_{2}$ is $0.75$. Policies $\pi_{3}$ and $\pi_{5}$, whose UCB are below this threshold, are identified as sub-optimal and consequently eliminated.
IV Performance Guarantee
The main theoretical results of this paper are the regret bound and the policy switching cost for our SDP-PE algorithm.
Theorem 6 (Performance guarantee of SDP-PE)